# In-Context Semi-Supervised Learning

### v1

@@ 1 Introduction
The Transformer (Attention_All) has revolutionized natural language processing, achieving remarkable success in tasks like language generation Attention_All; radford2019language; BERT; LLM_few; llama; deepseek; it has also been applied in many other areas, including vision ViT, network analysis cheng2025graph; graphTransformer_PEs; graphTransformer_PEs_rethink and biology Transformers_mol_biology. Motivated by the demonstrated effectiveness of Transformers for few-shot (or in-context) learning LLM_few, a promising line of recent research has focused on functional data vonoswald2023transformers; ahn2023transformers; Zhang23; Schlag21. In this setting one is given a set of contextual data pairs, where each pair is characterized by the observed input and output of a latent function. It has been demonstrated that a properly designed Transformer can in its forward pass perform functional gradient descent (GD) – or close variations thereof – for the function, and it can use that inference to predict the expected functional outcome for a specified query input to the function. The Transformer can perform this inference in-context, with no parameter fine-tuning.
Early work on Transformer-based in-context learning (ICL) focused on linear functions vonoswald2023transformers; ahn2023transformers; Zhang23; Schlag21, and that has been extended recently to nonlinear functions that reside in a reproducing kernel Hilbert space (RKHS) cheng2024transformers. These works focus on learning real-valued outcomes, and similar analysis was recently developed for categorical outcomes Aaron_ICL. Categorical observations are of interest when studying ICL for classification problems.
@@ 1.2 Related Work
Semi-supervised learning (SSL) has been a problem of long-standing interest to the machine learning community semi_sup_survey; deep_semi_sup; book_semi_sup. It is particularly valuable when labeled data are scarce, multiple unlabeled samples are observed at once, and there is an underlying structure associated with the covariates from both labeled and unlabeled data. Despite the recent surge in research on Transformer-based In-Context Learning (ICL) vonoswald2023transformers; ahn2023transformers; Zhang23; Schlag21; cheng2024transformers; Aaron_ICL, standard implementations often overlook the potential of the unlabeled portion of the context.
Recent work has begun to investigate this intersection, primarily focusing on scaling and data selection. agarwal2024many demonstrated that scaling to many-shot prompts yields significant gains and introduced Unsupervised ICL (using inputs only) to reduce dependence on curated labels. Similarly, chen2025maple improved many-shot ICL by influence-selecting unlabeled examples and pseudo-labeling them with an LLM. From a theoretical perspective, li2025and proved that while a single-layer linear-attention model cannot exploit unlabeled data, depth or looping constructs can implement polynomial estimators akin to the Expectation-Maximization (EM) algorithm, showing empirical gains on tabular tasks.
These findings motivate our focus on the fundamental mechanism by which Transformers leverage unlabeled context. Unlike previous approaches that rely on adding pseudo-labeled shots chen2025maple; agarwal2024many, we argue that Transformers can extract structure directly from unlabeled tokens. This relates to the field of spectral geometry and graph-based embeddings. Traditionally, embedding vectors for unlabeled samples—often derived from the eigenvectors of the graph Laplacian coifman2006diffusion—have been calculated offline and used as positional embeddings (PEs) in graph Transformers graphTransformer_PEs; graphTransformer_PEs_rethink. While cheng2025graph recently showed that Transformers can compute certain PEs during inference, we are unaware of any prior work that fully integrates this into a semi-supervised ICL framework.
@@ 2 Problem Setup
Background: In-Context Supervised Learning. For in-context supervised learning vonoswald2023transformers; cheng2024transformers, the Transformer is presented with contextual data ${\mathcal{C}}=\left\{(x^{(1)},y^{(1)}),\dots,(x^{(m)},y^{(m)})\right\}$, where $x^{(i)}\in\mathbb{R}^{d}$, and $y^{(i)}$ is a label. The query is $x^{(m+1)}$. The goal of prior work vonoswald2023transformers; cheng2024transformers was examination of the Transformer to infer $\mathbb{E}(y^{(m+1)})$ based on the contextual data. In this setting, it was assumed that there is an underlying function $f(x)$ responsible for the relationship between any input $x^{(i)}$ and output $y^{(i)}$. In most prior work of this type, investigation was performed on the potential for Transformers to infer context-dependent $f(x)$ on the forward pass, and use it to estimate $\mathbb{E}(y^{(m+1)})$.
In most prior work vonoswald2023transformers; cheng2024transformers the outcomes $y^{(i)}$ were real. Recently this has been extended to categorical $y^{(i)}\in\{1,\dots,C\}$ Aaron_ICL, and we will consider that setting here. Given $f(x)$, the relationship between $x^{(i)}$ and $y^{(i)}$ is modeled via the softmax (like at the output of language models Attention_All):
@@
For each category (label type) $c$ there is a learned embedding vector $w_{c}\in\mathbb{R}^{d^{\prime}}$, and the function $f(x)$ has a $d^{\prime}$-dimensional vector output. In-context learning (ICL) in this setting was first developed in Aaron_ICL, and for reader convenience details are provided in Appendix B.
@@ 3.1 Transformer for In-context Representation Learning
Given $\mathcal{C}$}, We define the affinity matrix $\mathcal{A}\in\mathbb{R}^{n\times n}$ with RBF bandwidth $h$ as $\mathcal{A}_{i,j}=\exp(-{\left\|x^{(i)}-x^{(j)}\right\|}_{2}^{2}/(2h))$, and $\mathcal{A}_{i,i}:=0$. Let $\mathcal{D}$ be the diagonal matrix defined as $\mathcal{D}_{i,i}=\sum_{j=1}^{n}\mathcal{A}_{ij}$. The discrete Laplacian $\operatorname{\mathcal{L}}:=\mathcal{D}-\mathcal{A}$, and its right normalized version $\hat{\operatorname{\mathcal{L}}}=I_{n\times n}-\mathcal{A}\mathcal{D}^{-1}$ are approximations to the Laplace-Beltrami operator coifman2006diffusion.
@@
Let $\tilde{\operatorname{\mathcal{L}}}=[{\sf TF}^{\operatorname{\mathcal{L}}}(Z_{0})]_{d:d+n}\in\mathbb{R}^{n\times n}$ denote the output of preceding Transformer stage. Given $\tilde{\operatorname{\mathcal{L}}}$ as input, the next stage ${\sf TF}^{\phi}$ takes as input $Z_{0}^{\top}=[\tilde{\operatorname{\mathcal{L}}}{{}^{\top}},0_{d\times d},I_{d\times k}]$. The architecture is a modification of the construction in cheng2024transformers for computing eigenvectors of graph Laplacians. In Lemma 2 provided in Appendix A.2, we show that there exists a Transformer ${\sf TF}^{\phi}$ that takes in $Z_{0}$ as defined above, and outputs the Laplacian Eigenmap for tokens 1n1...n. Concretely, let the output of ${\sf TF}^{\phi}$ be
@@ 3.2 Transformer for In-context Supervised Learning
Our derivation is based on the assumption that $f(\phi)$ resides in a reproducing kernel Hilbert space (RKHS) scholkopf2002learning, but the setup extends to softmax-based attention kernels as well Aaron_ICL; cheng2024transformers. From the RKHS perspective, let $f(\phi)=A\psi(\phi)$, with $\psi(\phi)$ a fixed mapping of features $\phi$ to a Hilbert space, and the parameters acting in that space are associated with matrix $A$.
@@
The form of the update in (6) is the same as that considered in previous Transformer ICL research vonoswald2023transformers; ahn2023transformers; Zhang23; Schlag21; cheng2024transformers; Aaron_ICL, and it naturally aligns with a self-attention layer (which considered real-valued observations).
@@ 4.2 ImageNet100: Semi-supervised IC-SSL on real data
We evaluate on ImageNet100, a 100-class subset of ILSVRC-2012 russakovsky2015imagenet commonly used for efficient representation learning caron2018deep; chen2020simple. Here, the “dataset size” denotes the number of in-context datasets (tasks). Each dataset has 200 batches, each batch sampling two random classes (50 images per class, 100 total) with replacement, and VGG-29 features simonyan2014vgg are used as inputs. The batch size is fixed at 200, with $3\%$ labeled samples (a $39\%$ variant is reported in Appendix F.4). For comparison, we use a scale-matched Transformer baseline (2-layer encoder, hidden dimension 128, $\sim$1.5M parameters), where class labels are projected into the same embedding space as the inputs and unlabeled samples are assigned a special unknown token, following the MetaICL formulation min2022metaicl to ensure fairness. Ground-truth EIG embeddings are obtained via Laplacian eigen-decomposition of the affinity graph as a spectral reference.
We compare three systems: (i) our constructed two‑stage model, (ii) a scale‑matched 2‑layer baseline Transformer trained end‑to‑end on the same episodes, and (iii) an ICL head on raw VGG features (Orig+ICL). We report standard ICL accuracy and a simple separation score caron2018deep; oord2018representation; CCI2024 (intra‑class similarity minus inter‑class similarity), which serves as a coarse proxy for representation quality (Figure 3).
@@
Table 1: Alignment metrics at labeled ratio $3\%$ and dataset size =5000=5000, comparing three embeddings: (Our model), (Transformer baseline), and (EIG). We report three alignment metrics (higher is better) introduced in the Platonic Representation Hypothesis tsuchiya2023platonic.
@@
Representational alignment. To connect these trends to geometry, we use mutual kkNN (mNN) alignment: for each sample, take its top-$k$ neighbors in two embeddings and compute the fraction of overlap between the two neighbor sets; the mNN score is the dataset-average (higher = more similar neighborhood structure) tsuchiya2023platonic. We compare three embeddings—ours, the baseline Transformer, and EIG (top-4 Laplacian eigenvectors from the affinity graph). Alignment rises with training data. Our model starts highly aligned to EIG (spectral inductive bias), while the baseline progressively aligns with both; with enough data, baseline–ours can exceed ours–EIG, and baseline–EIG improves but generally trails ours–EIG at moderate scales. Additional metrics (Cycle, LCS) show the same pattern; by $5000$ episodes all three pairs exhibit high alignment (Table 1).
@@ 4.3 Families of Manifolds and Their Geometry
Let $(\mathcal{M},g)$ denote a smooth, connected Riemannian manifold with intrinsic distance $d_{\mathcal{M}}$. We use five such manifolds, visualized in Figure  2. Each manifold admits a closed-form geodesic and a parametric map $\Phi\colon U\subset\mathbb{R}^{m}\!\to\!\mathbb{R}^{3}$. For example, for the sphere $\mathbb{S}^{2}$, $\Phi_{\mathrm{\mathbb{S}^{2}}}(\theta,\phi)=(\sin\theta\cos\phi,\;\sin\theta\sin\phi,\;\cos\theta)$, and its geodesic is $d_{\mathrm{\mathbb{S}^{2}}}(p,q)=\arccos(\langle p,q\rangle)$. Details of their parametric maps and geodesic derivations are in Appendix E. Many of these manifolds have had important applications in scientific machine learning (de2022riemannian; jing2022torsional).
@@ 4.4 Image Manifold Experiments
To test whether the representation-learning routine transfers to higher dimensional, natural-looking data, we build an image manifold with Stable Diffusion v1.5 podell2023sdxl. Concretely, we draw two latent noise vectors, treat them as endpoints, and generate a 100-step spherical linear interpolation in latent space. Each interpolated latent is decoded—without classifier guidance—into a $500\times 500$ RGB image. Repeating this over 500 prompt seeds yields 500 diverse sub-manifolds of images (each composed of 100 images). The intrinsic geometry of each sub-manifold is inherited from the latent interpolation; full generation and labeling details appear in Appendix E.3.
@@ A.1 Transformer for forming normalized graph Laplacian (Lemma 1)
The Transformer that we use for learning a Laplacian-based representation is based on the one proposed in Section 14.2 of cheng2025graph with one important modification – replacing the linear attention by RBF attention – enables the same architecture to efficiently compute the Laplacian of the nearest-neighbors graph in-context. See (9) above for definitions of $\operatorname{Attn}_{\operatorname{\texttt{rbf}}}$.
@@ A.2 Transformer for Computing Eigenmap (Lemma 2)
As with the preceding section, the Transformer that we use for computing the Eigenmap is based on the Transformer from in Lemma 9 of cheng2025graph with one important modification. Here, we retain the use of $\operatorname{Attn}_{\operatorname{\texttt{linear}}}$ as our goal is to implement a block-power-method algorithm for computing the Eigenvectors of a matrix. See (9) above for definitions of $\operatorname{Attn}_{\operatorname{\texttt{linear}}}$.
@@
Our proof is primarily based on Lemma 6, Corollary 7, and Lemma 9 of cheng2025graph. For ease of reference, we reproduce the necessary details below.
@@
Our proof proceeds in two steps. First, we verify that there exists a choice of parameters under which the Transformer orthogonalizes the $i^{th}$ column of $Z^{\Phi}_{\ell}$ with respect to all columns $j>i$. The construction is identical to Lemma 8 of cheng2025graph.

### v2

@@ 1 Introduction
The Transformer (Vaswani et al., 2017) has revolutionized natural language processing, achieving remarkable success in tasks like language generation Vaswani et al. (2017); Radford et al. (2019); Devlin et al. (2019); Brown et al. (2020); Touvron et al. (2023); DeepSeek (2025); it has also been applied in many other areas, including vision Dosovitskiy et al. (2021), network analysis Cheng et al. (2025); Kreuzer et al. (2021); Rampasek et al. (2022) and biology Jiang et al. (2024). Motivated by the demonstrated effectiveness of Transformers for few-shot (or in-context) learning Brown et al. (2020), a promising line of recent research has focused on functional data von Oswald et al. (2023); Ahn et al. (2023); Zhang et al. (2023); Schlag et al. (2021). In this setting one is given a set of contextual data pairs, where each pair is characterized by the observed input and output of a latent function. It has been demonstrated that a properly designed Transformer can in its forward pass perform functional gradient descent (GD) – or close variations thereof – for the function, and it can use that inference to predict the expected functional outcome for a specified query input to the function. The Transformer can perform this inference in-context, with no parameter fine-tuning.
Early work on Transformer-based in-context learning (ICL) focused on linear functions von Oswald et al. (2023); Ahn et al. (2023); Zhang et al. (2023); Schlag et al. (2021), and that has been extended recently to nonlinear functions that reside in a reproducing kernel Hilbert space (RKHS) Cheng et al. (2024). These works focus on learning real-valued outcomes, and similar analysis was recently developed for categorical outcomes Wang et al. (2025). Categorical observations are of interest when studying ICL for classification problems.
@@ 1.2 Related Work
Semi-supervised learning (SSL) has been a problem of long-standing interest to the machine learning community van Engelen and Hoos (2020); Ouali et al. (2020); Chapelle et al. (2010). It is particularly valuable when labeled data are scarce, multiple unlabeled samples are observed at once, and there is an underlying structure associated with the covariates from both labeled and unlabeled data. Despite the recent surge in research on Transformer-based In-Context Learning (ICL) von Oswald et al. (2023); Ahn et al. (2023); Zhang et al. (2023); Schlag et al. (2021); Cheng et al. (2024); Wang et al. (2025), standard implementations often overlook the potential of the unlabeled portion of the context.
Recent work has begun to investigate this intersection, primarily focusing on scaling and data selection. Agarwal et al. (2024) demonstrated that scaling to many-shot prompts yields significant gains and introduced Unsupervised ICL (using inputs only) to reduce dependence on curated labels. Similarly, Chen et al. (2025) improved many-shot ICL by influence-selecting unlabeled examples and pseudo-labeling them with an LLM. From a theoretical perspective, Li et al. (2025) proved that while a single-layer linear-attention model cannot exploit unlabeled data, depth or looping constructs can implement polynomial estimators akin to the Expectation-Maximization (EM) algorithm, showing empirical gains on tabular tasks.
These findings motivate our focus on the fundamental mechanism by which Transformers leverage unlabeled context. Unlike previous approaches that rely on adding pseudo-labeled shots Chen et al. (2025); Agarwal et al. (2024), we argue that Transformers can extract structure directly from unlabeled tokens. This relates to the field of spectral geometry and graph-based embeddings. Traditionally, embedding vectors for unlabeled samples—often derived from the eigenvectors of the graph Laplacian Coifman and Lafon (2006)—have been calculated offline and used as positional embeddings (PEs) in graph Transformers Kreuzer et al. (2021); Rampasek et al. (2022). While Cheng et al. (2025) recently showed that Transformers can compute certain PEs during inference, we are unaware of any prior work that fully integrates this into a semi-supervised ICL framework.
@@ 2 Problem Setup
Background: In-Context Supervised Learning. For in-context supervised learning von Oswald et al. (2023); Cheng et al. (2024), the Transformer is presented with contextual data ${\mathcal{C}}=\left\{(x^{(1)},y^{(1)}),\dots,(x^{(m)},y^{(m)})\right\}$, where $x^{(i)}\in\mathbb{R}^{d}$, and $y^{(i)}$ is a label. The query is $x^{(m+1)}$. The goal of prior work von Oswald et al. (2023); Cheng et al. (2024) was examination of the Transformer to infer $\mathbb{E}(y^{(m+1)})$ based on the contextual data. In this setting, it was assumed that there is an underlying function $f(x)$ responsible for the relationship between any input $x^{(i)}$ and output $y^{(i)}$. In most prior work of this type, investigation was performed on the potential for Transformers to infer context-dependent $f(x)$ on the forward pass, and use it to estimate $\mathbb{E}(y^{(m+1)})$.
In most prior work von Oswald et al. (2023); Cheng et al. (2024) the outcomes $y^{(i)}$ were real. Recently this has been extended to categorical $y^{(i)}\in\{1,\dots,C\}$ Wang et al. (2025), and we will consider that setting here. Given $f(x)$, the relationship between $x^{(i)}$ and $y^{(i)}$ is modeled via the softmax (like at the output of language models Vaswani et al. (2017)):
@@
For each category (label type) $c$ there is a learned embedding vector $w_{c}\in\mathbb{R}^{d^{\prime}}$, and the function $f(x)$ has a $d^{\prime}$-dimensional vector output. In-context learning (ICL) in this setting was first developed in Wang et al. (2025), and for reader convenience details are provided in Appendix B.
@@ 3.1 Transformer for In-context Representation Learning
Given $\mathcal{C}$}, We define the affinity matrix $\mathcal{A}\in\mathbb{R}^{n\times n}$ with RBF bandwidth $h$ as $\mathcal{A}_{i,j}=\exp(-{\left\|x^{(i)}-x^{(j)}\right\|}_{2}^{2}/(2h))$, and $\mathcal{A}_{i,i}:=0$. Let $\mathcal{D}$ be the diagonal matrix defined as $\mathcal{D}_{i,i}=\sum_{j=1}^{n}\mathcal{A}_{ij}$. The discrete Laplacian $\operatorname{\mathcal{L}}:=\mathcal{D}-\mathcal{A}$, and its right normalized version $\hat{\operatorname{\mathcal{L}}}=I_{n\times n}-\mathcal{A}\mathcal{D}^{-1}$ are approximations to the Laplace-Beltrami operator Coifman and Lafon (2006).
@@
Let $\tilde{\operatorname{\mathcal{L}}}=[{\sf TF}^{\operatorname{\mathcal{L}}}(Z_{0})]_{d:d+n}\in\mathbb{R}^{n\times n}$ denote the output of preceding Transformer stage. Given $\tilde{\operatorname{\mathcal{L}}}$ as input, the next stage ${\sf TF}^{\phi}$ takes as input $Z_{0}^{\top}=[\tilde{\operatorname{\mathcal{L}}}{{}^{\top}},0_{d\times d},I_{d\times k}]$. The architecture is a modification of the construction in Cheng et al. (2024) for computing eigenvectors of graph Laplacians. In Lemma 2 provided in Appendix A.2, we show that there exists a Transformer ${\sf TF}^{\phi}$ that takes in $Z_{0}$ as defined above, and outputs the Laplacian Eigenmap for tokens 1n1...n. Concretely, let the output of ${\sf TF}^{\phi}$ be
@@ 3.2 Transformer for In-context Supervised Learning
Our derivation is based on the assumption that $f(\phi)$ resides in a reproducing kernel Hilbert space (RKHS) Schölkopf and Smola (2002), but the setup extends to softmax-based attention kernels as well Wang et al. (2025); Cheng et al. (2024). From the RKHS perspective, let $f(\phi)=A\psi(\phi)$, with $\psi(\phi)$ a fixed mapping of features $\phi$ to a Hilbert space, and the parameters acting in that space are associated with matrix $A$.
@@
The form of the update in (6) is the same as that considered in previous Transformer ICL research von Oswald et al. (2023); Ahn et al. (2023); Zhang et al. (2023); Schlag et al. (2021); Cheng et al. (2024); Wang et al. (2025), and it naturally aligns with a self-attention layer (which considered real-valued observations).
@@ 4.2 ImageNet100: Semi-supervised IC-SSL on real data
We evaluate on ImageNet100, a 100-class subset of ILSVRC-2012 Russakovsky et al. (2015) commonly used for efficient representation learning Caron et al. (2018); Chen et al. (2020). Here, the “dataset size” denotes the number of in-context datasets (tasks). Each dataset has 200 batches, each batch sampling two random classes (50 images per class, 100 total) with replacement, and VGG-29 features Simonyan and Zisserman (2014) are used as inputs. The batch size is fixed at 200, with $3\%$ labeled samples (a $39\%$ variant is reported in Appendix F.4). For comparison, we use a scale-matched Transformer baseline (2-layer encoder, hidden dimension 128, $\sim$1.5M parameters), where class labels are projected into the same embedding space as the inputs and unlabeled samples are assigned a special unknown token, following the MetaICL formulation Min et al. (2022) to ensure fairness. Ground-truth EIG embeddings are obtained via Laplacian eigen-decomposition of the affinity graph as a spectral reference.
We compare three systems: (i) our constructed two‑stage model, (ii) a scale‑matched 2‑layer baseline Transformer trained end‑to‑end on the same episodes, and (iii) an ICL head on raw VGG features (Orig+ICL). We report standard ICL accuracy and a simple separation score Caron et al. (2018); Oord et al. (2018); Padhan and Sahu (2024) (intra‑class similarity minus inter‑class similarity), which serves as a coarse proxy for representation quality (Figure 3).
@@
Table 1: Alignment metrics at labeled ratio $3\%$ and dataset size =5000=5000, comparing three embeddings: (Our model), (Transformer baseline), and (EIG). We report three alignment metrics (higher is better) introduced in the Platonic Representation Hypothesis Tsuchiya et al. (2023).
@@
Representational alignment. To connect these trends to geometry, we use mutual kkNN (mNN) alignment: for each sample, take its top-$k$ neighbors in two embeddings and compute the fraction of overlap between the two neighbor sets; the mNN score is the dataset-average (higher = more similar neighborhood structure) Tsuchiya et al. (2023). We compare three embeddings—ours, the baseline Transformer, and EIG (top-4 Laplacian eigenvectors from the affinity graph). Alignment rises with training data. Our model starts highly aligned to EIG (spectral inductive bias), while the baseline progressively aligns with both; with enough data, baseline–ours can exceed ours–EIG, and baseline–EIG improves but generally trails ours–EIG at moderate scales. Additional metrics (Cycle, LCS) show the same pattern; by $5000$ episodes all three pairs exhibit high alignment (Table 1).
@@ 4.3 Families of Manifolds and Their Geometry
Let $(\mathcal{M},g)$ denote a smooth, connected Riemannian manifold with intrinsic distance $d_{\mathcal{M}}$. We use five such manifolds, visualized in Figure  2. Each manifold admits a closed-form geodesic and a parametric map $\Phi\colon U\subset\mathbb{R}^{m}\!\to\!\mathbb{R}^{3}$. For example, for the sphere $\mathbb{S}^{2}$, $\Phi_{\mathrm{\mathbb{S}^{2}}}(\theta,\phi)=(\sin\theta\cos\phi,\;\sin\theta\sin\phi,\;\cos\theta)$, and its geodesic is $d_{\mathrm{\mathbb{S}^{2}}}(p,q)=\arccos(\langle p,q\rangle)$. Details of their parametric maps and geodesic derivations are in Appendix E. Many of these manifolds have had important applications in scientific machine learning (De Bortoli et al., 2022; Jing et al., 2022).
@@ 4.4 Image Manifold Experiments
To test whether the representation-learning routine transfers to higher dimensional, natural-looking data, we build an image manifold with Stable Diffusion v1.5 Podell et al. (2023). Concretely, we draw two latent noise vectors, treat them as endpoints, and generate a 100-step spherical linear interpolation in latent space. Each interpolated latent is decoded—without classifier guidance—into a $500\times 500$ RGB image. Repeating this over 500 prompt seeds yields 500 diverse sub-manifolds of images (each composed of 100 images). The intrinsic geometry of each sub-manifold is inherited from the latent interpolation; full generation and labeling details appear in Appendix E.3.
@@ A.1 Transformer for forming normalized graph Laplacian (Lemma 1)
The Transformer that we use for learning a Laplacian-based representation is based on the one proposed in Section 14.2 of Cheng et al. (2025) with one important modification – replacing the linear attention by RBF attention – enables the same architecture to efficiently compute the Laplacian of the nearest-neighbors graph in-context. See (9) above for definitions of $\operatorname{Attn}_{\operatorname{\texttt{rbf}}}$.
@@ A.2 Transformer for Computing Eigenmap (Lemma 2)
As with the preceding section, the Transformer that we use for computing the Eigenmap is based on the Transformer from in Lemma 9 of Cheng et al. (2025) with one important modification. Here, we retain the use of $\operatorname{Attn}_{\operatorname{\texttt{linear}}}$ as our goal is to implement a block-power-method algorithm for computing the Eigenvectors of a matrix. See (9) above for definitions of $\operatorname{Attn}_{\operatorname{\texttt{linear}}}$.
@@
Our proof is primarily based on Lemma 6, Corollary 7, and Lemma 9 of Cheng et al. (2025). For ease of reference, we reproduce the necessary details below.
@@
Our proof proceeds in two steps. First, we verify that there exists a choice of parameters under which the Transformer orthogonalizes the $i^{th}$ column of $Z^{\Phi}_{\ell}$ with respect to all columns $j>i$. The construction is identical to Lemma 8 of Cheng et al. (2025).